OpenStudy (anonymous):

during a flu epidemic a company with 200 employees has 1/10 of their employees call in sick on monday and another 3/10 call in sick on tuesday. what is the total # of employees calling in sick during this 2 day period?

OpenStudy (anonymous):

1/10 of 200 = 1/10 x 200 = 200/10 = 20 people called in Monday 3/10 of 200 - 20 = 3/10 of 180 = 3/10 x 180 = 540/10 = 54 people called in Tuesday. I believe the answer is 54 + 20 = 74. You might need to get another opinion because I am not sure if the people that called in Monday were also out Tuesday.